### v2

@@ 5 Two-way Alternating Automata on Finite Words (2AFW)
6 From ALDLf Formulas to 2AFWs
Having defined the 2AFW, we are now ready to begin the first part of the ALDLf formula to NFW translation. The construction starts by converting from an ALDLf formula $\varphi$ in negation normal form to a a 2AFW $A_{\varphi}$ that defines the same language, i.e., $\mathcal{L}(\varphi)=\mathcal{L}(A_{\varphi})$. Definition 10 provides the formal definition. We provide an intuitive explanation here.
The states of $A_{\varphi}$ are the Fischer-Ladner closure of $\varphi$. The nodes of a run $\rho$ of $A_{\varphi}$ on a trace $\pi$ are labeled with states and positions (with the exception of the special case of leaf nodes labeled with configuration Accept; if a node $v$ of $\rho$ has state $\psi$ and position $i$ it is useful to think of the subtree with $v$ as root to be a run of $A_{\psi}$ on $\pi$ from $i$. Naturally, the start state of $A_{\varphi}$ is $\varphi$.
@@ 6 From ALDLf Formulas to 2AFWs
Rule 5 corresponds to rule A5. If a path automaton gets “stuck”, i.e., it is not in an accepting state and cannot make a transition, then all of the disjunctions evaluate to false. This is desirable because a stuck path automaton in this context represents that the first condition of rule A5 does not hold. Similarly, if the path automaton gets trapped in a loop and never reaches an accepting state, this also represents that the first condition of rule A5 does not hold. In this latter case, it results in a run with an infinite branch in which each node of the branch is labeled with a state of the form $\langle U\rangle\varphi$. Since there are no states of the form $[U]\varphi$, this is a rejecting branch, as desired. Since states of the form $\langle U\rangle\varphi$ “demand” reaching an appropriate position that satisfies $\varphi$, getting “stuck” in such a state means that the “demand” has not been met. Thus, states of the form $\langle U\rangle\varphi$ are not accepting states.
@@ 6.1 Construction
By Lemma 1 we have that $|CL(\varphi|\leq|\varphi|$. Now, $Q=CL(\varphi)$, so $|Q|\leq|\varphi|$. ∎
6.2 Example
@@
  2. 2.
    $\delta([\alpha_{3}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\{\})=([\alpha_{3}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},1)\wedge((\neg a,0)\vee[\alpha_{3_{s_{1}}}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},0)$
  3. 3.
@@
  4. 4.
    $\delta([\alpha_{3_{s_{1}}}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\{\})=(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},0)$
  5. 5.
    $\delta(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\{a\})=(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},-1)\vee(a,0)\wedge(\langle\alpha_{2_{q_{1}}}\rangle[\alpha_{1}]\textbf{false},0)$
  6. 6.
    $\delta(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\{\})=(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},-1)\vee(a,0)\wedge(\langle\alpha_{2_{q_{1}}}\rangle[\alpha_{1}]\textbf{false},0)$
  7. 7.
@@
  8. 8.
    $\delta(\langle\alpha_{2_{q_{1}}}\rangle[\alpha_{1}]\textbf{false},\{\})=([\alpha_{1}]\textbf{false},0)$
  9. 9.
@@
  10. 10.
    $\delta([\alpha_{1}]\textbf{false},\{\})=([\alpha_{1_{p_{1}}}]\textbf{false},-1)$
  11. 11.
@@
  12. 12.
    $\delta([\alpha_{1_{p_{1}}}]\textbf{false},\{\})=(\textbf{false},0)$
6.3 Correctness of ALDLf to 2AFW Construction
@@
Let $U=(R,T,\Delta,r,G)$, $\pi\in(2^{AP})^{+}$, $0\leq i<|\pi|$, $0\leq j<|\pi|$. Then $\pi,i,j\models U$ if and only if there is a vable path automaton walk of $U$ on $\pi$ from $i$ to $j$.
@@
7 From 2AFWs to NFWs
Now we proceed to the final step of the ALDLf to NFA translation: the construction of an NFA from a 2AFW. The technique we use here is a generalization of that given by Chandra, Kozen, and Stockmeyer to construct an NFA from a one-way alternating finite automaton [8].
@@ 7 From 2AFWs to NFWs
7.1 Definitions and Lemmas
7.1.1 History Strategy
@@
A history path $\beta$ in a history strategy $h$ on input $w$ is a maximal length sequence of pairs from $|w|\times Q$ that may be finite or infinite and has the following property: if $\beta=(u_{1},q_{1}),(u_{2},q_{2}),\ldots$ is infinite, then, for all $i\geq 0$, there is some $d_{i+1}\in\{-1,0,1\}$ such that $(d_{i+1},q_{i+1})\in h((u_{1},q_{1}),\ldots,(u_{i},q_{i})$ (if $\beta$ is finite, this condition holds for all $0\leq i<|\beta|$). Thus, $\beta$ is obtained by following transitions in the history strategy. We define $\inf(\beta)$ to be the set of states in $Q$ that occur infinitely often in $\beta$. We say that an infinite history path $\beta$ is accepting if there is a $q\in F$ such that $q\in\inf(\beta)$. We say that a finite history path $\beta$ of length $n$ is accepting if $\delta(q_{n},w_{u_{n}})=\textbf{true}$. We say that a history strategy $h$ is accepting if all of its history paths are accepting.
7.1.2 Strategy Word
@@ 8 Main Proofs
If condition 4 holds, then $i=|w|-1$ and $h(p)$ has an element of the form $(q,1)$. But this would imply that some node $x$ in $\rho$ has some child whose index is greater than $|w|-1$, which would mean that, per Definition 9, $\rho$ is not a run of $A$ on $w$. But $\rho$ is an accepting run of $A$ on $w$.
@@
Now, $G_{A,w}$ is a special instance of a parity game. A classic result for parity games is that if a parity game has a winning history strategy, then it has a memoryless winning strategy, i.e., it has a strategy that relies only on the current configuration of the game[9]. Let $h^{\prime}:(Q\times\mathbb{N})\rightarrow 2^{\{-1,0,1\}\times Q}$ be a winning memoryless strategy for $G_{A,w}$. Without loss of generality, $h^{\prime}$ is equivalent to the function $\gamma:\mathbb{N}\rightarrow 2^{Q\times\{-1,0,1\}\times Q}$. But $\gamma$ is an accepting strategy word over $w$.
@@
Let $A=(Q,\Sigma,\delta,q_{0},F)$ be a 2AFW and let $w$ be a word over the alphabet $\Sigma=2^{AP}$. There is an NFA $A_{n}$ such that $A_{n}$ accepts $w$ iff $A$ accepts $w$, $A_{n}$ has $2^{O(|Q|^{2})}$ states, and each state of $A_{n}$ is of size $O(|Q|^{4})$.
@@
To establish the lower bound, we note that the satisfiability problem for LTLf is PSPACE-complete and that a formula of LTLf can be converted to ALDLf in linear time and space using the following procedure: translate from LTLf to LDLf using the linear construction provided in [14], then translate from LDLf to ALDLf by converting each instance of path expression to a path automaton using Thompson’s construction[27].
@@
By Corollary 4.1, we have that there is an NFA $A_{n}$ such that $\mathcal{L}(A_{\varphi})=\mathcal{L}(A_{n})$ and that $A_{n}$ has $2^{O(|Q|^{2})}$ states and that each state is of size $O(|Q|^{4})$. Reachability from the initial state of $A_{n}$ to an accepting state of $A_{n}$ can be performed through a nondeterministic search: at each step, guess a letter of the alphabet and guess the successor state, while remembering only the successor state, and if an accepting state is reached as the final input is consumed then it is reachable from the initial state. Now, each letter is polynomial in the size of $\varphi$ and by Corollary 4.1 we have that each state of $A_{n}$ is polynomial in the size of $\varphi$. So this procedure can be done in NPSPACE. Since $\varphi$ and $A_{n}$ define the same language, $\varphi$ is satisfiable iff an accepting state of $A_{n}$ is reachable from the initial state of $A_{n}$. So satisfiability of $\varphi$ is in NPSPACE. But by Savitch’s Theorem [24], we have that NPSPACE = PSPACE. So satisfiability checking of $\varphi$ is in PSPACE.
9 Conclusion
@@
ALDLf extends the paradigm of LDLf by allowing the use of NFAs in place of regular expressions to express temporal constraints and also by providing for the direct expression of past modalities. Because NFAs are stateful, they can sometimes be more convenient to use than regular expressions. They are also exponentially more succinct than regular expressions. Satisfiability checking of ALDLf formulas is still PSPACE-complete, so these features come at no additional cost compared to LDLf.

### v4

@@ 5. Two-way Alternating Automata on Finite Words (2AFW)
6. From ALDLf Formulas to 2AFWs
Having defined the 2AFW, we are now ready to begin the first part of the ALDLf formula to NFW translation. The construction starts by converting from an ALDLf formula $\varphi$ in negation normal form to a 2AFW $A_{\varphi}$ that defines the same language, i.e., $\mathcal{L}(\varphi)=\mathcal{L}(A_{\varphi})$. Definition 10 provides the formal definition. We provide an intuitive explanation here.
The states of $A_{\varphi}$ are the Fischer-Ladner closure of $\varphi$. The nodes of a run $\rho$ of $A_{\varphi}$ on a trace $\pi$ are labeled with states and positions (with the exception of the special case of leaf nodes labeled with configuration Accept; if a node $v$ of $\rho$ has state $\psi$ and position $i$ it is useful to think of the subtree with $v$ as root to be a run of $A_{\psi}$ on $\pi$ from $i$.
Naturally, the start state of $A_{\varphi}$ is $\varphi$.
@@ 6. From ALDLf Formulas to 2AFWs
Rule 5 corresponds to rule A5. If a path automaton gets “stuck”, i.e., it is not in an accepting state and cannot make a transition, then all of the disjunctions evaluate to false. This is desirable because a stuck path automaton in this context represents that the first condition of rule A5 does not hold. Similarly, if the path automaton gets trapped in a loop and never reaches an accepting state, this also represents that the first condition of rule A5 does not hold. In this latter case, it results in a run with an infinite branch in which each node of the branch is labeled with a state of the form $\langle U\rangle\varphi$. Since there are no states of the form $[U]\varphi$, this is a rejecting branch, as desired.
Since states of the form $\langle U\rangle\varphi$ “demand” reaching an appropriate position that satisfies $\varphi$, getting “stuck” in such a state means that the “demand” has not been met. Thus, states of the form $\langle U\rangle\varphi$ are not accepting states.
@@ 6.1. Construction
By Lemma 1 we have that $|CL(\varphi|\leq|\varphi|$. Now, $Q=CL(\varphi)$, so $|Q|\leq|\varphi|$.
6.2. Example
@@
  2. (2)
    $\delta([\alpha_{3}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\emptyset)=([\alpha_{3}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},1)\wedge((\neg a,0)\vee[\alpha_{3_{s_{1}}}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},0)$
  3. (3)
@@
  4. (4)
    $\delta([\alpha_{3_{s_{1}}}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\emptyset)=(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},0)$
  5. (5)
    $\delta(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\{a\})=((\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},-1)\vee(a,0))\wedge((\langle\alpha_{2_{q_{1}}}\rangle[\alpha_{1}]\textbf{false},0))$
  6. (6)
    $\delta(\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},\emptyset)=((\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false},-1)\vee(a,0))\wedge((\langle\alpha_{2_{q_{1}}}\rangle[\alpha_{1}]\textbf{false},0))$
  7. (7)
@@
  8. (8)
    $\delta(\langle\alpha_{2_{q_{1}}}\rangle[\alpha_{1}]\textbf{false},\emptyset)=([\alpha_{1}]\textbf{false},0)$
  9. (9)
@@
  10. (10)
    $\delta([\alpha_{1}]\textbf{false},\emptyset)=([\alpha_{1_{p_{1}}}]\textbf{false},-1)$
  11. (11)
@@
  12. (12)
    $\delta([\alpha_{1_{p_{1}}}]\textbf{false},\emptyset)=(\textbf{false},0)$
6.3. Correctness of ALDLf to 2AFW Construction
@@
Let $U=(R,T,\Delta,r,G)$, $\pi\in(2^{AP})^{+}$, $0\leq i<|\pi|$, $0\leq j<|\pi|$. Then $\pi,i,j\models U$ if and only if there is a viable path automaton walk of $U$ on $\pi$ from $i$ to $j$.
@@
7. From 2AFWs to NFWs
Now we proceed to the final step of the ALDLf to NFA translation: the construction of an NFA from a 2AFW. The technique we use here is a generalization of that given by Chandra, Kozen, and Stockmeyer to construct an NFA from a one-way alternating finite automaton [CKSA81].
@@ 7. From 2AFWs to NFWs
7.1. Definitions and Lemmas
7.1.1. History Strategy
@@
A history path $\beta$ in a history strategy $h$ on input $w$ is a maximal length sequence of pairs from $|w|\times Q$ that may be finite or infinite and has the following property: if $\beta=(u_{1},q_{1}),(u_{2},q_{2}),\ldots$ is infinite, then, for all $i\geq 0$, there is some $d_{i+1}\in\{-1,0,1\}$ such that $(d_{i+1},q_{i+1})\in h((u_{1},q_{1}),\ldots,(u_{i},q_{i})$ (if $\beta$ is finite, this condition holds for all $0\leq i<|\beta|$).
Thus, $\beta$ is obtained by following transitions in the history strategy. We define $\inf(\beta)$ to be the set of states in $Q$ that occur infinitely often in $\beta$. We say that an infinite history path $\beta$ is accepting if there is a $q\in F$ such that $q\in\inf(\beta)$. We say that a finite history path $\beta$ of length $n$ is accepting if $\delta(q_{n},w_{u_{n}})=\textbf{true}$. We say that a history strategy $h$ is accepting if all of its history paths are accepting.
7.1.2. Strategy Word
@@ 8. Main Proofs
If condition 4 holds, then $i=|w|-1$ and $h(p)$ has an element of the form $(q,1)$.
But this would imply that some node $x$ in $\rho$ has some child whose index is greater than $|w|-1$, which would mean that, per Definition 9, $\rho$ is not a run of $A$ on $w$. But $\rho$ is an accepting run of $A$ on $w$.
@@
Now, $G_{A,w}$ is a special instance of a parity game. A classic result for parity games is that if a parity game has a winning history strategy, then it has a memoryless winning strategy, i.e., it has a strategy that relies only on the current configuration of the game [EJ91]. Let $h^{\prime}:(Q\times\mathbb{N})\rightarrow 2^{\{-1,0,1\}\times Q}$ be a winning memoryless strategy for $G_{A,w}$. Without loss of generality, $h^{\prime}$ is equivalent to the function $\gamma:\mathbb{N}\rightarrow 2^{Q\times\{-1,0,1\}\times Q}$. But $\gamma$ is an accepting strategy word over $w$.
@@
Let $A=(Q,\Sigma,\delta,q_{0},F)$ be a 2AFW and let $w$ be a word over the alphabet $\Sigma=2^{AP}$. There is an NFA $A_{n}$ such that $A_{n}$ accepts $w$ iff $A$ accepts $w$, $A_{n}$ has $2^{O(|Q|^{2})}$ states, and each state of $A_{n}$ is of size
$O(|Q|^{4})$.
@@
To establish the lower bound, we note that the satisfiability problem for LTLf is PSPACE-complete and that a formula of LTLf can be converted to ALDLf in linear time and space using the following procedure: translate from LTLf to LDLf using the linear construction provided in [DGV+13], then translate from LDLf to ALDLf by converting each instance of path expression to a path automaton using Thompson’s construction [Tho68].
@@
By Corollary 4.1, we have that there is an NFA $A_{n}$ such that $\mathcal{L}(A_{\varphi})=\mathcal{L}(A_{n})$ and that $A_{n}$ has $2^{O(|Q|^{2})}$ states and that each state is of size $O(|Q|^{4})$. Reachability from the initial state of $A_{n}$ to an accepting state of $A_{n}$ can be performed through a nondeterministic search: at each step, guess a letter of the alphabet and guess the successor state, while remembering only the successor state, and if an accepting state is reached as the final input is consumed then it is reachable from the initial state. Now, each letter is polynomial in the size of $\varphi$ and by Corollary 4.1 we have that each state of $A_{n}$ is polynomial in the size of $\varphi$. So this procedure can be done in NPSPACE. Since $\varphi$ and $A_{n}$ define the same language, $\varphi$ is satisfiable iff an accepting state of $A_{n}$ is reachable from the initial state of $A_{n}$. So satisfiability of $\varphi$ is in NPSPACE. But by Savitch’s Theorem [Sav70], we have that NPSPACE = PSPACE. So satisfiability checking of $\varphi$ is in PSPACE.
9. Conclusion
@@
ALDLf extends the paradigm of LDLf by allowing the use of NFAs in place of regular expressions to express temporal constraints and also by providing for the direct expression of past modalities, which can be combined with present and future modalities within the same formula. Because NFAs are stateful, they can sometimes be more convenient to use than regular expressions. They are also exponentially more succinct than regular expressions. Satisfiability checking of ALDLf formulas is still PSPACE-complete, so these features come at no additional cost compared to LDLf.